OpenStudy (anonymous):

Two 350.0-W speakers, A and B, are separated by a distance D = 3.59 m. The speakers emit in-phase sound waves at a frequency f = 10950.0 Hz. Point P1 is located at x1= 4.93 m and y1 = 0 m; point P2 is located at x2 = 4.93 m and y2 = –Δy. a) Neglecting speaker B, what is the intensity, IA1 (in W/m2), of the sound at point P1 due to speaker A? Assume that the sound from the speaker is emitted uniformly in all directions. b) What is this intensity in terms of decibels (sound level, βA1)? c) How far is P2 from P1, that is, what is Δy?

OpenStudy (anonymous):

... the intensity I of sound at any distance r from the source can be determined applying Inverse Square Law, \[I=\frac{W}{4\pi r^2}\] where \(W\) is the sound power in Watts ... this formula assume that your sound source emits sound wave equally in all directions... \(r\) is the distance of the listener from the sound source in meter...

OpenStudy (anonymous):

... the intensity in terms of dB, \[I(dB)=10\log \frac{I}{I_o}\]where \(I\) is the intensity in \(Watts/m^2\) and \(I_o=10^{-12} W/m^2\)

OpenStudy (anonymous):

...\(I_o\) is known as the Threshold of Hearing

OpenStudy (anonymous):

... for like sources of more than one, the total intensity in dB, \[I_{TOTAL}(dB)=I(dB)+10 \log n\] where \(n\) is the number of like sources...

OpenStudy (anonymous):

... hope this concepts help you...